2PH3  + 4O2 🡪  P2O5  +  3H2O


  1. How many grams of O2 will completely react with 49.0 g PH3?
1
Expert's answer
2021-10-17T11:19:32-0400

M(PH3) = 34.00 g/mol

n(PH3) "= \frac{49.0}{34.0} = 1.44 \;mol"

According to the reaction equation

n(O2) = 2n(PH3) "= 2 \times 1.44 = 2.88 \;mol"

M(O2) = 32.00 g/mol

m(O2) "= 2.88 \times 32 = 92.23 \;g"

Answer: 92.23 g